2. (10 points) For function f(x) = 4x3 – x4, find: (a) the critical points; (b) the open intervals on which the function is i please i need the answers asapp
Answer #1

4n3 – 4 f (se) = 12 x² - 4x3 f(s) = 4 x2 (3-2) a:) Critical points f(x) =0 4 se? (3-5) = 0 0,3 o 27 re- + > 0 0 3 b.) Incrc.) f(x) = 480² (3-2) f(x) = 248 - 12 80² = 122(2-2) f(x) = 0 at n=o,3 But at x=0, f(x) does not change from increasing to